OpenStudy (anonymous):

find all the zeros by factoring the function F(x)=X^3+8

OpenStudy (anonymous):

8=2^3 and we know that: a^3+b^3=(a+b)(a^2-ab+b^2)

OpenStudy (anonymous):

So, x^3 + 2^3 = ... can you expand it?

OpenStudy (anonymous):

x*x*x*2*2*2?

OpenStudy (anonymous):

a^3+b^3=(a+b)(a^2-ab+b^2) x^3+2^3=(x+2)(x^2-2x+4)

OpenStudy (anonymous):

see?

OpenStudy (anonymous):

yea. ok thank you.....again :)